OpenStudy (anonymous):

Determine the projection of u= 1.6i+3.3j in the direction v= -2.1i - 0.5j. a. 0.8i+0.2j b. 2.3i+0.5j c. -0.6i-0.3j d. -1.2i-1.2j

OpenStudy (anonymous):

I really absolutely do not get this, if you could help me solve it that would be much appreciated. Thank you thank you thank you.

OpenStudy (ksaimouli):

use the formula of projection on vectors

OpenStudy (anonymous):

I did that but I am getting so lost in my work its not even funny.

OpenStudy (ksaimouli):

\[u=<1.6,3.3> v<-2.1,-0.5>\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

and then i plug it into the formula so that its the dot vector multiplication over V squared with it all set to be multiplied by V again

OpenStudy (ksaimouli):

\[Proj=\frac{ a.b }{ |a|^2 }a\]

OpenStudy (anonymous):

yeah got that much then after plugging it all in...idek

OpenStudy (ksaimouli):

did u change to unit vector

OpenStudy (anonymous):

Nope. How do I do that?

OpenStudy (ksaimouli):

\[<a.b>=\frac{ a.b }{ \sqrt{a^2+b^2}} \]

OpenStudy (anonymous):

okay so multiply a times b all over the square of a squared + b squared, then after I get that what do I do with what I get from that

OpenStudy (anonymous):

wait, do i add the products of a times b?

OpenStudy (anonymous):

thank you for your help btw, i really appreciate it, math is NOT my subject

OpenStudy (ksaimouli):

\[\frac{ <-2.5,-0.5> }{ \sqrt{(-2.5)^2+(-0.5)^2} }\]

OpenStudy (anonymous):

okay then from there?

OpenStudy (ksaimouli):

show ur work where did you get

OpenStudy (anonymous):

Im solving right now but the one part of the equation is wrong, its -2.1 not -2.5.

OpenStudy (anonymous):

Okay so I have the top vector over 6.5

OpenStudy (anonymous):

then do I do that with the other one too?

OpenStudy (ksaimouli):

is the answer \[\frac{ -5.01 }{ 2.64}\]

OpenStudy (anonymous):

mmmm i didnt get that. Idek. Im just gonna drop this for now. Thank you for everything. You were helpful. I'm just math stupid